OpenStudy (shawn):

if f(x) = x^3 on [-2,1], is f odd?

OpenStudy (olivya999):

yes

OpenStudy (shawn):

how so ?

OpenStudy (math&ing001):

Take any -x with x from [-2,1] and test f(-x)

OpenStudy (shawn):

ok, take x = -2 f(-(-2)) = ??

OpenStudy (math&ing001):

Normally we work it with a variable to prove it for all x in the given domain. Like for x in [-2,1], f(-x)=(-x)^3=-x^3=-f(x) So f is odd on [-2,1].

OpenStudy (shawn):

so you're saying f(-(-2)) = -f(-2)? But 2 isn't in the domain

OpenStudy (math&ing001):

What exactly are trying to prove ?

OpenStudy (shawn):

if f is odd or not

OpenStudy (math&ing001):

So why do you go for specific values, you just need to prove it for that in the domain given. Fyi x^3 is odd for all real numbers i.e on \([-\infty,+\infty]\) ;]

OpenStudy (shawn):

oh ok. Thanks :)

OpenStudy (math&ing001):

Welcome =)